# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v3

@@ III Simultaneous symmetry breaking in many-body field theory
We proceed to review how symmetries emerge in many-body systems. For the sake of definiteness, we consider the following general time-independent second-quantization Hamiltonian, valid for both interacting bosons and fermions Fetter and Walecka (2003),
@@
$\hat{H}$ $=$ $\int\mathrm{d}\mathbf{x}~{}\hat{\Psi}^{\dagger}(\mathbf{x})\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V(\mathbf{x})\right]\hat{\Psi}(\mathbf{x})$
@@
where the field operator $\hat{\Psi}$, the external potential $V(\mathbf{x})$, and the interacting potential $\mathbf{V}(\mathbf{x}-\mathbf{x}^{\prime})$ may present a tensorial structure due to internal degrees of freedom (spin, pseudospin…). The corresponding Heisenberg equation of motion for the field operator is
$i\hbar\partial_{t}\hat{\Psi}(\mathbf{x},t)=[\hat{\Psi}(\mathbf{x},t),\hat{H}]=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V(\mathbf{x})\right]\hat{\Psi}(\mathbf{x},t)$
$+\int\mathrm{d}\mathbf{x}^{\prime}~{}\hat{\Psi}^{\dagger}(\mathbf{x}^{\prime},t)\mathbf{V}(\mathbf{x}-\mathbf{x}^{\prime})\hat{\Psi}(\mathbf{x}^{\prime},t)\hat{\Psi}(\mathbf{x},t).$ (9)
@@
$L\equiv\int\mathrm{d}\mathbf{x}\,\mathcal{L},~{}\mathcal{L}=i\hbar\Psi^{\dagger}\partial_{t}\Psi-\mathcal{H},$ (10)
with $\mathcal{H}$ the Hamiltonian density arising from Eq. (III). The canonical momentum associated to the field $\Psi$ is $\Pi(\mathbf{x})=i\hbar\Psi^{\dagger}(\mathbf{x})$, giving rise to the Poisson bracket
@@
This Lagrangian correspondence allows, via Noether’s theorem, to identify the conserved charges arising from the symmetries of the Hamiltonian. Specifically, any continuous symmetry transformation that leaves invariant the action gives rise to a conserved charge
@@
$x^{\prime\mu}=x^{\mu}+\alpha\xi^{\mu}(x)+\ldots,~{}\Psi^{\prime}(x)=\Psi(x)+\alpha\Delta\Psi(x)+\ldots$ (13)
@@
The most common symmetries are invariance under phase transformations, temporal and spatial translations, and rotations, i.e.,
$\Psi^{\prime}$ $=$ $\Psi e^{-i\theta},$ (14)
$t^{\prime}$ $=$ $t-t_{0},$
$\mathbf{x}^{\prime}$ $=$ $\mathbf{x}+\mathbf{x}_{0},$
$\mathbf{x}^{\prime}$ $=$ $R(\boldsymbol{\varphi})\mathbf{x},$
leading to the conserved charges
$Q_{\theta}$ $=$ $N=\int\mathrm{d}\mathbf{x}~{}\Psi^{\dagger}\Psi,$ (15)
$Q_{t}$ $=$ $E\equiv H=\int\mathrm{d}\mathbf{x}~{}\mathcal{H},$
$Q_{\mathbf{x}}$ $=$ $\mathbf{P}=\int\mathrm{d}\mathbf{x}~{}\Psi^{\dagger}(-i\hbar\nabla)\Psi,$
$Q_{\boldsymbol{\varphi}}$ $=$ $\mathbf{L}=\int\mathrm{d}\mathbf{x}~{}\Psi^{\dagger}[-i\hbar(\mathbf{x}\times\nabla)+\mathbf{S}]\Psi,$
which are the particle number, energy, momentum and angular momentum, respectively. In the latter case, the spin $\mathbf{S}$ arises due to the internal structure of the field $\Psi$.
In the following, lowercase Latin indices $a,b,c\ldots$ label both continuous symmetry transformations and the associated conserved charges, $a=\alpha,A$; lowercase Greek indices $\alpha,\beta,\gamma\ldots$ label symmetry transformations, $\alpha=\theta,t,\mathbf{x}\ldots$; and uppercase Latin indices $A,B,C\ldots$ label the conserved charges $Q_{\alpha,\beta,\gamma}$ associated to the symmetries $\alpha,\beta,\gamma$, $A=N,E,\mathbf{P}\ldots$ The continuous symmetry parameters, generators and conserved charges will be grouped in vectors $\boldsymbol{\alpha}$, $\mathbf{T}$, $\mathbf{Q}$, respectively.
@@
In particular, since most of the symmetries do not involve the time coordinate, $\xi^{0}=0$, and hence their conserved charges can be written as a simple bilinear expression in the field
$Q_{\alpha}=\hbar\int\mathrm{d}\mathbf{x}~{}\Psi^{\dagger}T_{\alpha}\Psi=\hbar\braket{\Psi}{T_{\alpha}}{\Psi},$ (17)
@@
$\braket{\chi}{\Psi}\equiv\int\mathrm{d}\mathbf{x}~{}\chi^{\dagger}(\mathbf{x})\Psi(\mathbf{x})$ (18)
the usual scalar product. The generators $T_{\alpha}$ are Hermitian operators under this scalar product, so symmetry transformations are implemented by unitary operators of the form $U=e^{-i\boldsymbol{\alpha}\cdot\mathbf{T}}$. For instance, for typical symmetries, $T_{\theta}=1$, $T_{\mathbf{x}}=-i\nabla$, $T_{\boldsymbol{\varphi}}=\mathbf{x}\times T_{\mathbf{x}}+\mathbf{S}$. We note that these derivations apply for an arbitrary many-body system, including both bosons and fermions.
III.1 Spontaneous symmetry breaking in the generalized Gibbs ensemble
We now address the case in which the state of the system spontaneously breaks $n$ symmetries, where in the following we restrict the vectors $\boldsymbol{\alpha},\mathbf{Q},\mathbf{T}$ to have $n$ components corresponding to those broken symmetries. Furthermore, we assume that those symmetries do not involve time [taking the form of Eq. (17)] and that their generators commute between themselves,
$[T_{\alpha},T_{\beta}]=0.$ (19)
@@ III.1 Spontaneous symmetry breaking in the generalized Gibbs ensemble
$L=\int\mathrm{d}\mathbf{x}~{}i\Psi^{*}\partial_{t}\Psi-H+\lambda^{\alpha}Q_{\alpha}.$ (20)
@@
$i\partial_{t}\Psi=\left[-\frac{\nabla^{2}}{2}+V(\mathbf{x})+|\Psi|^{2}\right]\Psi\equiv H_{\rm{GP}}\Psi,$ (22)
with $H_{\rm{GP}}$ the effective nonlinear GP Hamiltonian. This equation takes the same self-consistent form of Eq. (4) by choosing $X$ as a two-component vector $X=(\Psi,\Psi^{*})$.
@@
We remark that each component $T_{\alpha}\Psi_{0}$ of the vector $\mathbf{T}\Psi_{0}$ is nonvanishing as they are broken symmetries by assumption. Furthermore, we also assume in the following that they are linearly independent, meaning that there is no constant vector $\boldsymbol{u}$ satisfying $u^{\alpha}T_{\alpha}\Psi_{0}=0$. If this were the case, the conserved charges would not be independent between themselves because then $u^{\alpha}Q_{\alpha}=0$. Nevertheless, even in this situation, one could further restrict the vectors $\boldsymbol{\alpha},\mathbf{T},\mathbf{Q}$ in such a way that the components of $\mathbf{T}\Psi_{0}$ are all linearly independent.
The standard time-independent GP equation is retrieved from the GPG equation for $\lambda^{\alpha}=0,~{}\alpha\neq\theta$, and $\lambda^{\theta}=\mu$, where $\mu$ is the chemical potential, the Lagrange multiplier associated to the particle number $N$, emerging due to the spontaneous symmetry-breaking of the $U(1)$-invariance under phase transformations. When additional continuous symmetries are broken, the generalized Gibbs ensemble allows us to treat all of them on equal grounds, leading to the GPG equation.
The actual dynamics of the symmetry-breaking state is obtained by inserting the stationary GPG wavefunction in the original time-dependent GP equation (22). Taking as initial condition $\Psi(\mathbf{x},0)=\Psi_{0}(\mathbf{x})$ gives
$i\partial_{t}\Psi=H_{GP}\Psi=(\boldsymbol{\lambda}\cdot\mathbf{T})\Psi\Longrightarrow\Psi(\mathbf{x},t)=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\Psi_{0}(\mathbf{x}).$ (24)
@@
$=$ $\int\mathrm{d}\mathbf{x}\left[\Psi_{0}^{*}H_{GP}\Psi_{0}-\frac{|\Psi_{0}|^{4}}{2}\right]=\lambda^{\alpha}Q_{\alpha}-\int\mathrm{d}\mathbf{x}\,\frac{|\Psi_{0}|^{4}}{2}.$
@@
$Q_{\beta}\partial_{A}\lambda^{\beta}=\int\mathrm{d}\mathbf{x}\,\frac{\partial_{A}|\Psi_{0}|^{4}}{2},$ (27)
@@
$\partial_{A}E=\frac{\partial E}{\partial Q_{\alpha}}=\lambda^{\alpha}\equiv\lambda_{A}$ (28)
where, for any vector $v$ with upper indices $\alpha$, we lower indices as $v_{A}\equiv v^{\alpha}$. This notation will be justified in Sec. III.2, from where a simpler derivation of this result can be obtained [see Eq. (48) and ensuing discussion]:
$\partial_{A}E$ $=$ $\int\mathrm{d}\mathbf{x}~{}(\partial_{A}\Psi_{0}^{*})H_{GP}\Psi_{0}+(H_{GP}\Psi_{0})^{*}\partial_{A}\Psi_{0}$ (29)
$=$ $i\lambda^{\beta}(z_{A}|z_{\beta})=\lambda^{\alpha}.$
Interestingly, since $p=|\Psi_{0}|^{4}/2$ is the local pressure, we can rewrite Eq. (III.1) as
$E=\lambda^{\alpha}Q_{\alpha}-\int\mathrm{d}\mathbf{x}~{}p,$ (30)
@@
$\int\mathrm{d}V~{}\partial_{A}p=Q_{\beta}\partial_{A}\lambda^{\beta}.$ (32)
III.2 Nambu-Goldstone modes in the generalized Gibbs ensemble
@@
$i\partial_{t}\Phi=M(t)\Phi,\,\Phi=\left[\begin{array}[]{c}\varphi\\ \varphi^{*}\end{array}\right],\,M(t)=\left[\begin{array}[]{cc}N(t)&A(t)\\ -A^{*}(t)&-N^{*}(t)\end{array}\right],$ (33)
@@
$N(t)=-\dfrac{\nabla^{2}}{2}+V(\mathbf{x})+2|\Psi(\mathbf{x},t)|^{2},~{}A(t)=\Psi(\mathbf{x},t)^{2}.$ (34)
@@
$i\partial_{t}z_{a}=M(t)z_{a},~{}z_{a}\equiv\left[\begin{array}[]{l}\partial_{a}\Psi\\ \partial_{a}\Psi^{*}\end{array}\right].$ (35)
@@
$M_{0}z_{\alpha}=0,~{}M_{0}=\left[\begin{array}[]{cc}N_{0}&A_{0}\\ -A_{0}^{*}&-N_{0}^{*}\end{array}\right],$ (37)
@@
$N_{0}=-\dfrac{\nabla^{2}}{2}+V(\mathbf{x})+2|\Psi_{0}(\mathbf{x})|^{2}-\boldsymbol{\lambda}\cdot\mathbf{T},~{}A_{0}=\Psi^{2}_{0}(\mathbf{x}).$ (38)
We can then identify $z_{\alpha}$ as the NG mode associated to the spontaneous symmetry breaking in a particular application of the general result of Eq. (7). In turn, each NG mode $z_{\alpha}$ is paired with a Gibbs mode $z_{A}$, obtained by taking derivative of the GPG equation with respect to the corresponding conserved charge $Q_{\alpha}$ [see Eq. (26)],
@@
We note that the linear independence of the NG modes $z_{\alpha}$ is already guaranteed by the assumption of the linear independence between the components of the vector $\mathbf{T}\Psi_{0}$ [see discussion after Eq. (23)]. As an example, consider a one-dimensional (1D) plane wave solution $\Psi_{0}(x)=e^{iqx}$ in a homogeneous system, which apparently spontaneously breaks both $U(1)$ and spatial-translation symmetry. However, since $T_{x}\Psi_{0}=qT_{\theta}\Psi_{0}=q\Psi_{0}$, then $z_{x}=qz_{\theta}$, and we only have one independent NG mode. Indeed, the resulting BdG dispersion relation is that of a homogeneous system plus a Doppler shift Recati et al. (2009). The momentum is also proportional to the particle number as $P=qN$. In general, when the system is in the ground state, there are counting rules relating the number of independent NG modes and broken-symmetry generators Watanabe and Murayama (2012); Hidaka (2013); Watanabe (2020).
@@
The matrix $M_{0}$ is a time-independent linear operator, so the solutions to Eq. (42) can be always decomposed in terms of a complete set of eigenmodes,
$M_{0}z_{n}=\omega_{n}z_{n},~{}z_{n}\equiv\left[\begin{array}[]{c}u_{n}\\ v_{n}\end{array}\right].$ (43)
These modes form an orthonormal basis under the inner product
$(z_{n}|z_{m})\equiv\braket{z_{n}}{\sigma_{z}z_{m}}=\int\mathrm{d}\mathbf{x}~{}u^{*}_{n}u_{m}-v^{*}_{n}v_{m},$ (44)
since $M_{0}$ is pseudo-Hermitian, $(z_{n}|M_{0}z_{m})=(M_{0}z_{n}|z_{m})$, where $\braket{z_{n}}{z_{m}}$ is the usual scalar product for two spinors and $\sigma_{i}$ are the Pauli matrices. The pseudo-Hermiticity of $M_{0}$ also implies that the inner product is conserved for two solutions of the time-dependent BdGG equation (42). However, both $M_{0}$ and the inner product are not positive definite. Indeed, by defining conjugate modes as $\bar{z}_{n}\equiv\sigma_{x}z_{n}^{*}$, it is easy to see that
$M_{0}\bar{z}_{n}=-\omega^{*}_{n}\bar{z}_{n},~{}~{}\,(z_{n}|z_{m})=-(\bar{z}_{m}|\bar{z}_{n})=-(\bar{z}_{n}|\bar{z}_{m})^{*}.$ (45)
In the following, for simplicity, we assume that there are no dynamical instabilities (i.e., there are no complex frequency modes) and that the only modes with zero frequency are the NG modes. It is immediate to show that one can separate the BdGG modes into two orthogonal sectors: the Goldstone-Gibbs sector formed by the paired modes $\{z_{a}\}=\{z_{\alpha},z_{A}\}$, and the regular Bogoliubov sector formed by the finite-frequency eigenmodes $\{z_{n}\}$,

### v4

@@ 3 Simultaneous symmetry breaking in stationary states
We proceed to review how symmetries emerge in many-body systems. For the sake of definiteness, we consider the following general time-independent second-quantization Hamiltonian, valid for both interacting bosons and fermions [73],
@@
$\hat{H}=\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ \hat{\Psi}^{\dagger}(\mathbf{x})\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V(\mathbf{x})\right]\hat{\Psi}(\mathbf{x})$ (8)
@@
where the field operator $\hat{\Psi}$, the external potential $V(\mathbf{x})$, and the interacting potential $\mathbf{V}(\mathbf{x}-\mathbf{x}^{\prime})$ may present a tensorial structure due to internal degrees of freedom (spin, pseudospin…). The corresponding Heisenberg equation of motion for the field operator $\hat{\Psi}(\mathbf{x},t)=e^{i\hat{H}t/\hbar}\hat{\Psi}(\mathbf{x})e^{-i\hat{H}t/\hbar}$ is
$i\hbar\partial_{t}\hat{\Psi}(\mathbf{x},t)=[\hat{\Psi}(\mathbf{x},t),\hat{H}$]
$=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V(\mathbf{x})\right]\hat{\Psi}(\mathbf{x},t)$
$+\int\mathrm{d}\mathbf{x}^{\prime}\leavevmode\nobreak\ \hat{\Psi}^{\dagger}(\mathbf{x}^{\prime},t)\mathbf{V}(\mathbf{x}-\mathbf{x}^{\prime})\hat{\Psi}(\mathbf{x}^{\prime},t)\hat{\Psi}(\mathbf{x},t).$ (9)
@@
$L\equiv\int\mathrm{d}\mathbf{x}\,\mathcal{L},\leavevmode\nobreak\ \mathcal{L}=i\hbar\Psi^{\dagger}\partial_{t}\Psi-\mathcal{H},$ (10)
with $\mathcal{H}$ the Hamiltonian density arising from Eq. (8). The canonical momentum associated to the field $\Psi$ is $\Pi(\mathbf{x})=i\hbar\Psi^{\dagger}(\mathbf{x})$, giving rise to the Poisson bracket
@@
This Lagrangian correspondence allows, via Noether theorem, to identify the conserved charges arising from the symmetries of the Hamiltonian. Specifically, any continuous symmetry transformation that leaves invariant the action gives rise to a conserved charge
@@
$x^{\prime\mu}$ $=$ $x^{\mu}+\alpha\xi^{\mu}(x)+\ldots$
$\Psi^{\prime}(x)$ $=$ $\Psi(x)+\alpha\Delta\Psi(x)+\ldots$ (13)
@@
The most common symmetries are invariance under phase transformations, temporal and spatial translations, and rotations, leading to the conservation of particle number $N$, energy $E$ and momentum $\mathbf{P}$, and angular momentum $\mathbf{L}$:
$\Psi^{\prime}=\Psi e^{-i\theta}$ $\Longrightarrow N=\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ \Psi^{\dagger}\Psi,$ (14)
$t^{\prime}=t-t_{0}$ $\Longrightarrow E\equiv H=\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ \mathcal{H},$
$\mathbf{x}^{\prime}=\mathbf{x}+\mathbf{x}_{0}$ $\Longrightarrow\mathbf{P}=\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ \Psi^{\dagger}(-i\hbar\boldsymbol{\nabla})\Psi,$
$\mathbf{x}^{\prime}=R(\boldsymbol{\varphi})\mathbf{x}$ $\Longrightarrow\mathbf{L}=\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ \Psi^{\dagger}[-i\hbar(\mathbf{x}\times\boldsymbol{\nabla})+\mathbf{S}]\Psi,$
where, in the last line, the spin $\mathbf{S}$ arises due to the internal structure of the field $\Psi$.
In the following, lowercase Greek indices $\alpha,\beta\ldots$ label symmetry transformations, $\alpha=\theta,t\ldots$, while uppercase Latin indices $A,B,\ldots$ label the conserved charges $Q_{\alpha,\beta,\ldots}=N,E,\ldots$ associated to the symmetries $\alpha,\beta,\ldots$ The continuous symmetry parameters, the generators and the conserved charges are grouped in vectors $\boldsymbol{\alpha}$, $\mathbf{T}$, $\mathbf{Q}$, respectively. Finally, lowercase Latin indices $a,b,\ldots$ indistinctively label continuous symmetry transformations or the associated conserved charges, $a=\alpha,A$. A summary of the notation is presented in Table 1.
@@
In particular, most of the symmetries do not involve the time coordinate, $\xi^{0}=0$, and hence their conserved charges can be written as a simple bilinear expression in the field
$Q_{\alpha}=\hbar\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ \Psi^{\dagger}T_{\alpha}\Psi=\hbar\braket{\Psi|T_{\alpha}|\Psi},$ (16)
@@
$\braket{\chi|\Psi}\equiv\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ \chi^{\dagger}(\mathbf{x})\Psi(\mathbf{x})$ (17)
the usual scalar product. The generators $T_{\alpha}$ are Hermitian operators under this scalar product, so symmetry transformations are implemented by unitary operators of the form $U=e^{-i\boldsymbol{\alpha}\cdot\mathbf{T}}$. We note that these derivations apply for an arbitrary many-body system, including both bosons and fermions.
3.1 Spontaneous symmetry breaking in the generalized Gibbs ensemble
We now address the case in which the state of the system spontaneously breaks $n$ continuous symmetries, where in the following we restrict the vectors $\boldsymbol{\alpha},\mathbf{Q},\mathbf{T}$ to have $n$ components corresponding to those broken symmetries. Furthermore, we assume that those symmetries do not involve time, so their charges take the form of Eq. (16), and that their generators commute between themselves,
$T_{\alpha},T_{\beta}]=0.$ (18)
@@ 3.1 Spontaneous symmetry breaking in the generalized Gibbs ensemble
$L=\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ i\Psi^{\dagger}\partial_{t}\Psi-H+\lambda^{\alpha}Q_{\alpha}.$ (19)
@@
Symmetry Generator Charge
Label $\alpha$ $T_{\alpha}$ $A\equiv Q_{\alpha}$
Vector $\boldsymbol{\alpha}$ $\mathbf{T}$ $\mathbf{Q}$
Example $(\theta,t,\mathbf{x})$ $(1,i\partial_{t},-i\boldsymbol{\nabla})$ $(N,E,\mathbf{P})$
Table 1: Summary of notation employed in this work.
@@
$i\partial_{t}\Psi=\left[-\frac{\nabla^{2}}{2}+V(\mathbf{x})+|\Psi|^{2}\right]\Psi\equiv H_{GP}\Psi,$ (21)
with $H_{GP}$ the effective nonlinear GP Hamiltonian. This equation takes the same self-consistent form of Eq. (4) by choosing $\mathbf{X}$ as a two-component vector $\mathbf{X}=[\Psi,\Psi^{*}]^{T}$, so then $H(\mathbf{X})$ is a diagonal matrix, $H(\mathbf{X})=\textrm{diag}[H_{GP},-H_{GP}]$.
@@
We remark that each component $T_{\alpha}\Psi_{0}$ of the vector $\mathbf{T}\Psi_{0}$ is nonvanishing as they reflect broken symmetries by assumption. Furthermore, we also assume in the following that they are linearly independent, meaning that there is no constant vector $\boldsymbol{u}$ satisfying $u^{\alpha}T_{\alpha}\Psi_{0}=0$. If this were the case, the conserved charges would not be independent between themselves because then $u^{\alpha}Q_{\alpha}=0$. Nevertheless, even in this situation, one could further restrict the vectors $\boldsymbol{\alpha},\mathbf{T},\mathbf{Q}$ in such a way that the components of $\mathbf{T}\Psi_{0}$ are all linearly independent.
The standard time-independent GP equation is retrieved from the GPG equation for $\lambda^{\alpha}=0,\leavevmode\nobreak\ \alpha\neq\theta$, and $\lambda^{\theta}=\mu$, where $\mu$ is the chemical potential, the Lagrange multiplier associated to the particle number $N$, emerging due to the spontaneous symmetry breaking of the $U(1)$-invariance under phase transformations. When additional continuous symmetries are spontaneously broken, the generalized Gibbs ensemble allows us to treat all of them on equal grounds, leading to the GPG equation.
The actual dynamics of the symmetry-breaking state is obtained by inserting the stationary GPG wavefunction into the original time-dependent GP equation (21). Taking as initial condition $\Psi(\mathbf{x},0)=\Psi_{0}(\mathbf{x})$ gives
$i\partial_{t}\Psi=H_{GP}\Psi=(\boldsymbol{\lambda}\cdot\mathbf{T})\Psi\Longrightarrow$ (23)
$\Psi(\mathbf{x},t)=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\Psi_{0}(\mathbf{x}).$
@@
$=$ $\int\mathrm{d}\mathbf{x}\left[\Psi_{0}^{*}H_{GP}\Psi_{0}-\frac{|\Psi_{0}|^{4}}{2}\right]$
$=$ $\lambda^{\alpha}Q_{\alpha}-\int\mathrm{d}\mathbf{x}\,\frac{|\Psi_{0}|^{4}}{2}.$
@@
$\int\mathrm{d}\mathbf{x}\,\frac{\partial_{A}|\Psi_{0}|^{4}}{2}=Q_{\beta}\partial_{A}\lambda^{\beta},$ (26)
@@
$\partial_{A}E=\frac{\partial E}{\partial Q_{\alpha}}=\lambda^{\alpha}\equiv\lambda_{A},$ (27)
where, for any vector $v$ with upper indices $\alpha$, we lower indices as $v_{A}\equiv v^{\alpha}$. This notation will be justified in Sec. 3.3, from where a simpler derivation of this result can be obtained [see Eq. (51) and ensuing discussion]:
$\partial_{A}E$ $=\int\mathrm{d}\mathbf{x}\,\left[(\partial_{A}\Psi_{0}^{*})H_{GP}\Psi_{0}+(H_{GP}\Psi_{0})^{*}\partial_{A}\Psi_{0}\right]$
$=i\lambda^{\beta}(z_{A}|z_{\beta})=\lambda^{\alpha}.$ (28)
Interestingly, since $p=|\Psi_{0}|^{4}/2$ is the local pressure, we can rewrite Eq. (3.1) as
$E=\lambda^{\alpha}Q_{\alpha}-\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ p,$ (29)
@@
$\int\mathrm{d}V\leavevmode\nobreak\ \partial_{A}p=Q_{\beta}\partial_{A}\lambda^{\beta}.$ (31)
Since the continuous symmetry parameters and their associated charges are conjugate position/momentum variables in a Hamiltonian language, these thermodynamic relations explain the role of the Lagrange multipliers as generalized velocities,
$\dot{\alpha}=\partial_{A}H=\lambda^{\alpha}.$ (32)
On the other hand, since the Hamiltonian is invariant under the symmetry transformations themselves, we immediately retrieve charge conservation, $\dot{Q}_{\alpha}=-\partial_{\alpha}H=0$.
3.2 Nambu-Goldstone modes in the generalized Gibbs ensemble
@@
$i\partial_{t}\Phi=M(t)\Phi,\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi=\left[\begin{array}[]{c}\varphi\\ \varphi^{*}\end{array}\right],$ (33)
@@
$M(t)$ $=$ $\left[\begin{array}[]{cc}N(t)&A(t)\\ -A^{*}(t)&-N^{*}(t)\end{array}\right],$ (36)
$N(t)$ $=$ $-\dfrac{\nabla^{2}}{2}+V(\mathbf{x})+2|\Psi(\mathbf{x},t)|^{2},$ (37)
$A(t)$ $=$ $\Psi(\mathbf{x},t)^{2}.$
@@
$i\partial_{t}z_{a}=M(t)z_{a},\leavevmode\nobreak\ z_{a}\equiv\left[\begin{array}[]{l}\partial_{a}\Psi\\ \partial_{a}\Psi^{*}\end{array}\right].$ (38)
@@
$M_{0}z_{\alpha}=0,\leavevmode\nobreak\ M_{0}=\left[\begin{array}[]{cc}N_{0}&A_{0}\\ -A_{0}^{*}&-N_{0}^{*}\end{array}\right],$ (40)
@@
$N_{0}=-\dfrac{\nabla^{2}}{2}+V(\mathbf{x})+2|\Psi_{0}(\mathbf{x})|^{2}-\boldsymbol{\lambda}\cdot\mathbf{T},\leavevmode\nobreak\ A_{0}=\Psi^{2}_{0}(\mathbf{x}).$ (41)
We can then identify $z_{\alpha}$ as the NG mode associated to the spontaneous symmetry breaking in a particular application of the general result of Eq. (7). In turn, each NG mode $z_{\alpha}$ is paired with a Gibbs mode $z_{A}$, obtained by taking derivative in the GPG equation with respect to the corresponding conserved charge $Q_{\alpha}$ [see Eq. (25)],
@@
We note that the linear independence of the NG modes $z_{\alpha}$ is already guaranteed by the assumption of the linear independence between the components of the vector $\mathbf{T}\Psi_{0}$ [see discussion after Eq. (22)]. As an example, consider a one-dimensional (1D) plane wave solution $\Psi_{0}(x)=e^{iqx}$ in a homogeneous system, which apparently spontaneously breaks both $U(1)$ and spatial-translation symmetry. However, since $T_{x}\Psi_{0}=qT_{\theta}\Psi_{0}=q\Psi_{0}$, then $z_{x}=qz_{\theta}$, and we only have one independent NG mode. Indeed, the resulting BdG dispersion relation is that of a homogeneous system plus a Doppler shift [74]. The momentum is also proportional to the particle number as $P=qN$. In general, when the system is in the ground state, there are counting rules relating the number of independent NG modes and broken-symmetry generators [75, 76, 77].
@@
The matrix $M_{0}$ is a time-independent linear operator, presenting eigenmodes of the form
$M_{0}z_{n}=\epsilon_{n}z_{n},\leavevmode\nobreak\ z_{n}\equiv\left[\begin{array}[]{c}u_{n}\\ v_{n}\end{array}\right].$ (46)
These modes are orthogonal under the inner product
$(z_{n}|z_{m})\equiv\braket{z_{n}|\sigma_{z}z_{m}}=\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ [u^{*}_{n}u_{m}-v^{*}_{n}v_{m}],$ (47)
since $M_{0}$ is pseudo-Hermitian, $(z_{n}|M_{0}z_{m})=(M_{0}z_{n}|z_{m})$, where $\braket{z_{n}|z_{m}}$ is the usual scalar product for two spinors and $\sigma_{i}$ are the Pauli matrices. The pseudo-Hermiticity of $M_{0}$ also implies that the inner product between two solutions of the time-dependent BdGG equation (45) is conserved. However, both $M_{0}$ and the inner product are not positive definite. Indeed, by defining conjugate modes as $\bar{z}_{n}\equiv\sigma_{x}z_{n}^{*}$, it is easy to see that
$M_{0}\bar{z}_{n}$ $=$ $-\epsilon^{*}_{n}\bar{z}_{n},$ (48)
$(z_{n}|z_{m})$ $=$ $-(\bar{z}_{m}|\bar{z}_{n})=-(\bar{z}_{n}|\bar{z}_{m})^{*}.$
In the following, for simplicity, we assume that there are no dynamical instabilities (i.e., there are no complex eigenvalues) and that the only zero modes are the NG modes. It is immediate to show that one can separate the BdGG modes into two orthogonal sectors: the Goldstone-Gibbs sector formed by the paired modes $\{z_{a}\}=\{z_{\alpha},z_{A}\}$, and the regular Bogoliubov sector formed by the eigenmodes $\{z_{n}\}$,